OpenStudy (anonymous):

What is the value of z in the equation 5z − 9 = 36? 4 5 6 9

Vocaloid (vocaloid):

add 9 to both sides first, and then tell me what you get

OpenStudy (anonymous):

wait wht

Vocaloid (vocaloid):

left side is 5x - 9 right side is 36 add 9 to the left side, add 9 to the right side, and then tell me what you get

OpenStudy (jadeishere):

You need to isolate the variable, in order to do that, you take -9 away from each sides, so to make 0 you add -9 + 9, then add 9 to the right side of the equation because whatever you do to one side, you do to the other >.<

OpenStudy (anonymous):

I had got 14 and 45

OpenStudy (jadeishere):

Okay, so when you add 9 to 36, you get 45 so now you have 5z = 45 in order to COMPLETELY isolate the variable you need to divide both sides by 5 because of \(5\) (z) so now you have z = 45/5

OpenStudy (anonymous):

\[5z-9=36\\5z=36+9\\5z=45\\z=\frac{ 45 }{ 9 }\\z=5\ ans\]

OpenStudy (jadeishere):

Um, wrong @enione

OpenStudy (anonymous):

Just plug in the numbers to z and see if it equals to the answer

OpenStudy (jadeishere):

\(45/5 = 9\)

OpenStudy (anonymous):

oh! no, \[z= 36\div 5\\z=9 Ans\]

OpenStudy (anonymous):

thers no 9 for my answer choice tho

OpenStudy (jadeishere):

Yesh there is

OpenStudy (anonymous):

lol nvm I seen it know I was looking at a different problem its they same way

OpenStudy (anonymous):

ha! ha! i'm drunk.